A quantity, its 2/3, its 1/2 ,and its 1/7 , added together, become 9. What is the quantity?

A) 42/97 B)197/42 C) 873/42 D)378/55


If possible please show me how to do a problem like this.

Answer by venugopalramana(3286)

A quantity, its 2/3, its 1/2 ,and its 1/7 , added together, become 9. What is the quantity?
ASSUME THE QUANTITY AS X..AND PROCEED AS IF X IS A KNOWN NUMBER LIKE 5 OR 4 OR 10...OR SOME NUMBER.
A quantity, its 2/3 = X*2/3=2X/3
its 1/2 = X*/2=X/2
,and its 1/7 =X*1/7=X/7
, added together..WE GET...2X/3 + X/2 + +X/7 = (2*14+ 1*21 +1*6 )X/42=55X/42= 9
X=42*9/55=378/55..IS THE ANSWER